OpenStudy (anonymous):

can someone please explain how to do this briefly. For each of the sequence of numbers:  Identify the type of sequence.  Determine the next two terms of the sequences.  Describe a rule which can be used to form the sequence. (a) 81; 108; 144; : : : (b) 60, 80, 320/3 ; : : :

OpenStudy (amistre64):

identify; is it arithmatic, or geometric what expression can generate the terms what are the next 2 terms

OpenStudy (amistre64):

if 108-81 = 144-108, then its arithmatic if 108/81 = 144/108; it geometris

OpenStudy (amistre64):

216 not= 225 so its not a simple arith. might be off by a constant

OpenStudy (amistre64):

no, still should work out .... might have to see a pattern

OpenStudy (amistre64):

120 240 320 ... you sure there aint no typos?

OpenStudy (amistre64):

180 + 60 ----- 240 +60 ---- 300 not=320

OpenStudy (amistre64):

144 -108 ----- 36 not= 81